A class of 24 students have an average test score of 92. A second class averaged 82. If both classes are combined, the average on the test is 88. How many students are in the second class?
Answer by stanbon(75887)

A class of 24 students have an average test score of 92. A second class averaged 82. If both classes are combined, the average on the test is 88. How many students are in the second class?
-------
1st class has 24*92 = 2208 points
2nd class has 82x points (x is the # of students in that class)
----
The combined classes have 88(24+x) points
--------
Equation:
2208 + 82x = 88(24+x)
2208 + 82x = 88*24 + 88x
6x = 2208-88*24
6x = 96
x = 16 (# of students in the 2nd class)
-------------------
Cheers,
Stan H.
===============